OpenStudy (anonymous):

this is an implicit differentiation... (xy)^1/2=x+3y Please help :)

OpenStudy (anonymous):

( √u ) ' = ...? u = xy --> u' = ....?

OpenStudy (anonymous):

@lizlozada How much you can fill in?

OpenStudy (anonymous):

ummm 1/2(xy)^-1/2(xy'+y)+1+3y'

OpenStudy (anonymous):

*=1+3y

OpenStudy (anonymous):

I want you to fill in what I'm asking!

OpenStudy (anonymous):

sorry that's the best i can do because we didn't learn with u, we learned like that

hartnn (hartnn):

what u have written is correct! do u need to isolate y' ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats my struggle

hartnn (hartnn):

let me give u hint from the question : (xy)^(1/2) = x+3y so (xy)^(-1/2) = 1/ (xy)^(1/2) = ?

hartnn (hartnn):

and right side is 1+3y'

OpenStudy (anonymous):

what happened to the 1/2 in front of (xy)^(-1/2)

hartnn (hartnn):

it will be there, 1/2 (xy)^(-1/2) = 1/2 (xy)^(1/2) = ? do u need to prove y' = something, or just find y'

OpenStudy (anonymous):

just find it

hartnn (hartnn):

then forget that hint. just distribute [1/2(xy)^-1/2](xy'+y)=1+3y' xy' [1/2(xy)^-1/2] + y [1/2(xy)^-1/2] = 1+3y' now collect y' terms on one side

hartnn (hartnn):

got that ? i just distributed [1/2(xy)^-1/2]

OpenStudy (anonymous):

distributed into what?

hartnn (hartnn):



hartnn (hartnn):